# A Simple and Effective $L_{2}$ Norm-Based Strategy for KV Cache Compression

### v1

@@ 4 Related Works
Recently, various long-context LLMs, such as Gemini-Pro-1.5 (Reid et al., 2024), Claude-3 (Anthropic, 2024), and GPT4 (Achiam et al., 2023), have shown the promising capability to process hundred thousands of tokens in the context. The increased number of input lengths results in a high decoding latency; thus, there has been a growing interest in speeding up the decoding with long contexts. Some works propose efficient memory management strategies to reduce the IO time overheads, e.g., PageAttention (Kwon et al., 2023), Infinite-LLM (Lin et al., 2024) and vAttention (Prabhu et al., 2024). Another line of research focuses on compressing the KV cache to improve efficiency. DMC (Nawrot et al., 2024) compresses KV cache by dynamically merging tokens while requiring expensive continual pre-training. For fine-tuning free compression strategy, H2O (Zhang et al., 2024) identifies important KV pairs by leveraging the attention scores from all queries, FastGen (Ge et al., 2023b) leverages the different attention patterns in different heads for compression, and SnapKV (Li et al., 2024) selects KV pairs based on attention scores from user’s query. Unlike these works, our method only utilises the $L_{2}$ norm of embedding for compression without leveraging the attention information, and to the best of our knowledge, we are the first to find that the influence of a KV pair can be determined by $L_{2}$ norm. Previous work (Darcet et al., 2024) finds the hidden states with high $L_{2}$ norm usually aggregate more important and global information. On the other hand, our findings indicate that a low $L_{2}$ norm of key embedding generally results in a high attention score.
5 Conclusion
@@
In this paper, we introduced a simple yet highly effective strategy for KV cache compression in LLMs based on the $L_{2}$ norm of key embeddings. We show that there is a significant correlation between the $L_{2}$ norm of a key embedding and its attention score. Leveraging this observation, our proposed method compresses the KV cache by retaining only those keys with the lowest $L_{2}$ norm. Our experimental results on various tasks show that our compression strategy maintains the predictive accuracy of the model while significantly reducing the memory footprint. Our approach is straightforward and can be applied directly to any transformer-based, decoder-only LLM.
@@ 6 Limitations
While our research provides valuable insights, our experiments and analyses were conducted exclusively using the Llama2-7b, Llama2-7b-32k and Llama2-7b-80k models. In future work, we will test on Llama2-13b and Llama2-70b, as well as the Llama3 family, to verify that our findings generalise to other models and architectures. Additionally, while we show that the $L_{2}$ norm played a significant role in our experiments, we do not have a comprehensive theoretical explanation for why this is the case. Understanding the underlying reasons behind the importance of the $L_{2}$ norm would require further theoretical exploration and empirical validation.
Appendix A Further Analysis
We discuss further the correlation between $L_{2}$ norm and attention scores. In Fig. 7, we show the ALr from Eq. 1 for two heads from layer 10 and layer 0. We see that in layer 0, the difference is larger due to a lower correlation.
In Fig. 8, we display $L_{2}$ norm and attention correlation across heads and layers using the original Llama2-7b and the long context Llama2-7b-32k and Llama2-7b-80k. We can see that patterns are quite consistent across all the models.
(a) Layer-7 Head-10, which has a high correlation between attention score and $L_{2}$-Norm.
(b) Layer-0 Head-0, which has a low correlation between attention score and $L_{2}$-Norm.
Figure 7: Attention loss of ideal compression and $L_{2}$ norm-based compression in Llama-2-7b-80k. The $x$-axis represents the compression ratio; the $y$-axis represents the attention loss (defined by Eq. 1) The results average over 1024 chunks on Wikipedia, with a length of 1024.
(a) Llama-2-7b.
(b) Llama-2-7b-80k
(c) Llama-2-7b-longlora-32k-ft
Figure 8: $\mathcal{Y}$ (defined by Eq. 3) of different models. The light square means a high difference between $L_{2}$-Norm. We observe that llama-2 and its fine-tuned long-context versions have a consistent pattern – the heads in the first two layers and middle layers are more different than others.
Appendix B More results on Language modelling task
@@
Figure 12: Skipping compression at different layers
Appendix C More Results on Long-Context Modelling Tasks
In addition to llama-2-7b-80k (Fu et al., 2024), we test the compression method using llama-2-7b-longlora-32k-ft (Chen et al., 2023) on the needle-in-a-haystack and passkey retrieval tasks. As shown in Fig. 13(a), we can see that compressing 30% of KV cache only results in a slight performance degradation on the needle-in-a-haystack task. We also observe that the performance even increases slightly when we compress 10% of KV cache. In figure Fig. 13(b), we observe that the llama-2-7b-longlora-32k-ft maintains 100% performance when compressing 80% of KV cache and only as a slight decrease when compressing 90% of KV cache. Furthermore, the model fails to generate correct answers if we compress KV pairs with low $L_{2}$ norm and keep high $L_{2}$ norm ones. The evaluation results of llama-2-7b-longlora-32k-ft are consistent with the llama-2-7b-80k, which further indicates the effectiveness of compressing KV cache using $L_{2}$ norm.
C.1 Analysis of Skipped Layers
@@
Appendix D Experimental setup
Appendix E More Visualizations

### v2

@@ 6 Related Work
Recently, various long-context LLMs, such as Gemini-Pro-1.5 (Reid et al., 2024), Claude-3 (Anthropic, 2024), and GPT4 (Achiam et al., 2023), have shown the promising capability to process hundred thousands of tokens in the context. The increased number of input lengths results in a high decoding latency; thus, there has been a growing interest in speeding up the decoding with long contexts. Some works propose efficient memory management strategies to reduce the IO time overheads, e.g., PageAttention (Kwon et al., 2023), Infinite-LLM (Lin et al., 2024) and vAttention (Prabhu et al., 2024). Another line of research focuses on compressing the KV Cache to improve efficiency. DMC (Nawrot et al., 2024) compresses KV Cache by dynamically merging tokens while requiring expensive continual pre-training. For fine-tuning free compression strategy, H2O (Zhang et al., 2024b) identifies important KV pairs by leveraging the attention scores from all queries, FastGen (Ge et al., 2023a) leverages the different attention patterns in different heads for compression, and SnapKV (Li et al., 2024) selects KV pairs based on attention scores from user’s query. Unlike these works, our method only utilises the $L_{2}$ norm of embedding for compression without leveraging the attention information, and to the best of our knowledge, we are the first to find that the influence of a KV pair can be determined by $L_{2}$ norm. Previous work (Darcet et al., 2024) finds the hidden states with high $L_{2}$ norm usually aggregate more important and global information. On the other hand, our findings indicate that a low $L_{2}$ norm of key embedding generally results in a high attention score.
7 Conclusions
@@
In this paper, we introduced a simple yet highly effective strategy for KV Cache compression in LLMs based on the $L_{2}$ norm of key embeddings. We show that there is a significant correlation between the $L_{2}$ norm of a key embedding and its attention score. Leveraging this observation, we compress the KV Cache by retaining only those keys with the lowest $L_{2}$ norm. Our experimental results on various tasks show that our compression strategy maintains the predictive accuracy of the model while significantly reducing the memory footprint. Our approach is straightforward and can be applied directly to any transformer-based, decoder-only LLM.
@@ 8 Limitations
While our research offers valuable insights, we tested only on relatively small models (Llama family and Gemma up to 8 billion parameters). In future work, we will assess our method on larger-scale models to ensure our findings generalize Additionally, while we show that the $L_{2}$ norm played a significant role in our experiments, we do not have a comprehensive theoretical explanation for why this is the case. Understanding the underlying reasons behind the importance of the $L_{2}$ norm would require further theoretical exploration and empirical validation. Finally, we observed (Figure 2) that compressing based on $L_{2}$ norm can be less effective depending on the layer and head considered, and we intend to investigate per-head compression ratios to leverage this observation.
Appendix A More results on Language modelling task
@@
Figure 13: Skipping compression at different layers with Llama2-7b
Appendix B More Results on Long-Context Modelling Tasks
In addition to llama-2-7b-80k [Fu et al., 2024], we test the compression method using llama-2-7b-longlora-32k-ft [Chen et al., 2023] on the needle-in-a-haystack and passkey retrieval tasks. As shown in Fig. 15(a), we can see that compressing 30% of KV Cache only results in a slight performance degradation on the needle-in-a-haystack task. We also observe that the performance even increases slightly when we compress 10% of KV Cache. In figure Fig. 15(b), we observe that the llama-2-7b-longlora-32k-ft maintains 100% performance when compressing 80% of KV Cache and only as a slight decrease when compressing 90% of KV Cache. Furthermore, the model fails to generate correct answers if we compress KV pairs with low $L_{2}$ norm and keep high $L_{2}$ norm ones. The evaluation results of llama-2-7b-longlora-32k-ft are consistent with the llama-2-7b-80k, which further indicates the effectiveness of compressing KV Cache using $L_{2}$ norm.
B.1 Analysis of Skipped Layers
@@
B.2 Longbench Evaluation
In this section we show detailed results from the LongBench dataset [Zhang et al., 2024a]. In Figure 21 we show results for Llama2-80k, while in Figure 22 we show results for the long context model Llama3.1-8b.
Figure 21: Evaluation results of Llama-2-7b-80k on long context tasks from Longbench, including narrativeqa and qasper, hotpotqa, 2wikimqa, and qmsum.
Figure 22: Evaluation results of Llama-3.1-8B on long context tasks from Longbench, including narrativeqa and qasper, hotpotqa, 2wikimqa, and qmsum.
Appendix C More Visualizations
@@
Appendix D Additional token embeddings plots
We show in Figure 29 some additional figure that represent Llama3-8b token embeddings sparsity.
Figure 29: Key projections of Llama3-8b of the bos $|beginoftext|$ token vs other tokens. Each value represents the activation in a specific dimension for the embedding of the key projection. We found similar patterns across almost all heads and layers and in multiple texts.
Appendix E Experimental setup